solve by addition method: 3/7x + 5/9y = 27

1/9 x + 2/7y = 7

Answer by stanbon(75887)

1st: 3/7x + 5/9y = 27
2nd: 1/9 x + 2/7y = 7
-------------------
Multiply 1st by 63 and 2nd by 63 to get rid of the denominators:
3rd: 27x + 35y = 27*63
4th: 7x + 18y = 7*63
-------------
Multiply 3rd by 7 and 4th by 27 to get:
5th: 189x + 245y = 189*63
6th: 189x + 486y = 189*63
--------------
Subtract 5th from 6th to get:
241y = 0
y = 0
------------
Substitute into (3/7)x+(5/9)y = 27 to solve for "x":
(3/7)x = 27
x = (7/3)27 = 63
========================================
Cheers,
Stan H.